OpenStudy (anonymous):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? (Points : 4) 9, 40, 42 5, 12, 13 6, 9, 10 13, 35, 37

OpenStudy (welshfella):

see if they fit the pythagoras theorem c^2 = a^2 + b^2 where c is the hypotenuse ( the longest side ) of the triangle

OpenStudy (welshfella):

for example lets look at no 3:- longest side = 10 10^2 = 100 9^2 = 81 6^2 = 36 add 117 100 is not equal to 117 so 3 is not a right angled triangle

OpenStudy (welshfella):

do you get the idea?

OpenStudy (anonymous):

sorry this alot hard for me. can you tell me witch is it